OpenStudy (anonymous):

Salmon often jump waterfalls to reach their breeding grounds. One salmon starts 2.00m from a waterfall that is 0.55m tall and jumps at an angle of 32.0 degrees. What must be the salmon's minimum speed to reach the waterfall?

OpenStudy (abb0t):

\[\theta = 32.0º \] \[\Delta d_v = 0.55 m \] \[\Delta d_h = 2.0 m \] \[g = -9.8 \frac{ m }{ s^2 } \] Trying to find: \[V_i\] \[V_f = 0 \frac{ m}{s}\] \[V_f^2=V_i^2+2g \Delta d_v \] \[0=V_i^2+2g \Delta d_v \] rearrange using algebra to find: \[V_i\]